# Leveraging Structural Knowledge in Diffusion Models for Source Localization in Data-Limited Graph Scenarios

### v2

@@ 5.4. Analysis of Pretraining on Simulation Data
To investigate how pretraining on simulation data affects model’s few-shot/zero-shot performance under different data volume, we evaluate SIDSL’s F1 scores across pretraining dataset with different propagation model combinations (IC, LT, and IC+LT with ratio 1:1) and compare it with non-pretrained (NP) SIDSL. We then test them under different ratios of the finetune dataset of real-world propagation data. The results of four datasets are shown in Figure 4.
Finetune data volume analysis. Across all four datasets, models with pretraining consistently outperformed non-pretrained (NP) models in low-data scenarios. Under IC+LT combination, the pretrained models achieved optimal F1 scores on Digg and Twitter using just 5% of training data, while NP models required 50%. Similarly, on Android and Christianity datasets, pretrained models reached near-peak performance (within 0.033 and 0.018 F1) with only 20% of the data, compared to NP models which needed 50%. These findings suggest that aligning pretraining patterns with specific propagation characteristics benefits model performance.
Pattern analysis. On the Digg dataset, SIDSL pretrained with IC+LT and IC configurations outperform that using LT, indicating that IC patterns better align with Digg’s propagation characteristics, as also shown in (Zhang et al., 2024). For Android and Christianity Q&A datasets, IC+LT and LT configurations performed better in few-shot scenarios. This is likely because user participation in these platforms’ information cascades (e.g., comments) is driven by cumulative influence from multiple sources (LT) rather than single independent triggers (IC). On Twitter, IC+LT configuration performs the best, demonstrating that pattern diversity enhances pretraining effectiveness on this dataset.
@@ 5.5. Ablation Study
We conduct ablation studies to investigate the significance of key components in SIDSL. For the first ablation model, SIDSL w/o SBD (structure-prior biased diffusion), instead of computing estimation through Equation 1, we directly substitute it with the node having the highest closeness centrality in the infection subgraph. For the second ablation, SIDSL w/o GNN-LP, we replace the observation state encoding with a standalone GNN while maintaining an equivalent parameter count.
Table 4 presents the ablation results on Digg and Christianity datasets under both full (100%) and limited (10%) training data conditions. Removing either component leads to decreased overall performance in both 100% and 10% data conditions. There are two more findings: First, removing SBD leads to more severe performance degradation compared to removing GNN-LP, with the impact most pronounced in recall metrics. The substantial decline in recall (TP/(TP+FN)) indicates a reduction in true positive predictions (as TP+FN remains constant for the same test set). Notably, the decline in recall exceeds that of precision (TP/(TP+FP)), which also implies a decrease in positive predictions (TP+FP). This pattern suggests that removing SBD causes the model to predict fewer positive samples overall, demonstrating SBD’s role in maintaining the model’s predictive balance. Second, the impact of removing GNN-LP is more pronounced under limited data conditions, with F1 scores dropping by -17% to -21% at 10% data versus -4% to -7% at full data. This demonstrates GNN-LP’s particular effectiveness in low-resource scenarios.
@@ 5.6. Time Cost Analysis
In inference time, our model is the second highest because of the iterative nature of the DDPM-based denoising process. We opt for DDPM as our foundation due to its classical design and proven effectiveness. It’s worth highlighting that SIDSL’s architecture is fully compatible with more computationally efficient diffusion variants, such as DDIM (Song et al., 2020), which could substantially reduce the current computational overhead. This flexibility, combined with our model’s transfer capabilities, makes SIDSL particularly resource-efficient in practical deployments.
@@
5.7. Results on Synthetic Propagation Patterns
@@
It is a common practice in previous works (Ling et al., 2022; Yan et al., 2024; Wang et al., 2022) to evaluate methods’ performance on synthetic data with established propagation patterns. Following (Ling et al., 2022; Wang et al., 2022), we conducted experiments on two commonly evaluated networks: Net Science (Net) and Power Grid(Power). Using both IC and LT models, we simulated 100 steps until convergence to generate two sets of synthetic data. Strong baseline models were trained and tested on these datasets, with results shown in Figure 6. Our method outperforms the baselines, achieving F1 score improvements of 7.6% and 4.7% under the IC model, and 2.7% and 1.7% under the LT model on the two datasets respectively. These results demonstrate our method’s ability to accurately identify sources with established propagation patterns.
6. Conclusion
In this paper, we presented SIDSL, a structure-prior informed diffusion framework for robust source localization in graphs with limited real-world propagation data. By incorporating topology-aware priors through graph label propagation, a GNN-parameterized label propagation module, and a structure-prior biased denoising process, our framework effectively addresses three key challenges: unknown propagation patterns, complex topology-propagation relationships, and class imbalance issues. Extensive experiments on four real-world datasets demonstrate SIDSL’s superior performance, achieving 7.5-13.3% improvements in F1 scores over state-of-the-art methods. More importantly, through effective pretraining with synthetic data, SIDSL maintains robust performance with only 10% training data, showing remarkable advantages in both few-shot and zero-shot learning scenarios. These results validate our framework’s strong generalization capability and practical applicability in real-world source localization tasks where labeled data is scarce. Future directions include improving model efficiency through hierarchical decomposition approaches to handle larger-scale networks while maintaining localization accuracy.
Appendix A Comparison of multiple source localization methods
In Table 5, we compare the functionality, requirements, and application scenarios of mainstream source localization methods. “Ind.” refers to whether the method considers modeling the indeterminacy of source locations. ”Applicable patterns” refers to the specific propagation pattern to which the method can be applied. ”Obs. input” refers to the required input for the method to detect the sources. ”PI” refers to whether the data-driven method is informed by prior knowledge.
From the demonstration, we can observe that our method is designed to be the most functional and capable of handling a broader range of real-world applications. Our method also requires less input data, which is more practical. The method proposed holds significant practical value and addresses the limitations of the existing methods. Additionally, as another method based on the diffusion model, DDMSL and TGASI require the propagation process data during training and the acquisition or calculation of parameters for the infectious model before source localization. This limitation restricts the model’s practical application value. Also, PGSL resembles SLVAE’s framework and merely utilizes a flow-based model to replace the VAE in SLVAE, while our diffusion model exhibits stronger distribution modeling capabilities. GINSD considers incomplete user data scenarios and utilizes a positional embedding module to distinguish incomplete nodes in the source inference process, and as we do not consider such circumstances, GINSD reduces to a simple GAT-based baseline similar to GCNSI.
It should also be noted that two recent works (Wang et al., 2024; Ling et al., 2024) focus on source localization in a cross-platform setting, which is orthogonal to our research problem and thus not discussed.
Appendix B Analyzing source centrality in empirical data
Figure 7. (a)(b) Normalized closeness centrality of infected users and sources for all cascades in Digg. The blue histogram shows the normalized closeness centrality distribution of infected nodes, while the red one shows that of source nodes. (a) is the density distribution of closeness centrality. The dashed line indicates the mean centrality for each node type. (b) is the frequency distribution of closeness centrality. The orange box highlights the part where centrality is above 0.8. (c) The closeness centrality(CC) probability density function of the predicted and ground truth source nodes on the Digg dataset. The blue histogram shows the normalized closeness centrality distribution of the ground truth sources, while the red one shows that of the predicted ones.
We believe that the source centrality assumption is not only common in most existing propagation patterns and real-world scenarios, as evidenced by the literature (Ali et al., 2020; Dong et al., 2019), but also validated by the competitive performance on real-world datasets of baselines like LPSI and GCNSI, which are devised based on similar assumptions. We show the analytical results demonstrating the effectiveness of the assumption in the following.
In our analysis of the real-world dataset Digg, we evaluate the normalized(max-min) closeness centrality density and frequency of the source nodes in the subgraph consisting of infected nodes to partially reflect the centrality characteristic of the sources. The closeness centrality (CC) specifically reflects the node topological distance to all other nodes in the subgraph, rather than its degree attribute. The result is shown in Figure 7(a). From Figure 7(a), the mean normalized closeness centrality of sources is higher than the average of all infected nodes, and source nodes cover over 63% of the nodes with the centrality score exceeding 0.8, as shown in Figure 7(b). The overall results demonstrate the crucial role of source nodes in the information diffusion process and their higher likelihood of being central to the network structure within the cascades.
Our proposed method, rather than strictly adhering to the centrality in outputting prediction results, exhibits stronger expressive capabilities. Intuitively, on homogeneous networks—where the probability of propagation along network edges is the same and fixed—the assumption can be strictly applied to locate the source of propagation. Such a propagation pattern that strictly obeys the centrality assumption is an indispensable subset that can be covered by the propagation patterns our model can characterize. As the proposed model leverages a simulated label propagation conditional module based on the centrality assumption but employs a graph neural network to learn the influence of the network’s heterogeneous topology from the data, other circumstances can also be modeled when learning from the data within our flexible data-driven framework. We have statistically analyzed the closeness centrality (CC) probability density function of the source nodes predicted by our trained SIDSL model on the Digg dataset and compared it with the ground truth centrality of the source nodes in Figure 7(c). The mean and standard deviation for the CCs of the predicted sources are 0.7020 and 0.1444, and that for the CCs of the ground truth sources are 0.7044 and 0.1567, showing that there is no harmful bias in our method’s prediction. This statistical result indicates that our model captures the source distribution observed in empirical data, not just theoretical derivations. Our method not only uses knowledge to guide inference to accelerate learning but also learns distribution patterns beyond the knowledge, from the data.
Figure 8. Additional experiments for simulated SIR scenarios on basic performance comparison.
Appendix C Additional Results of Performance under Other Propagation Patterns
We conduct additional experiments for simulated SIR scenarios on basic performance comparison. The results are shown in Figure 8. It can be seen that our model can still achieve competitive results compared to these baselines, proving our method’s applicability. The results also indicate that in terms of precision, ours achieved the highest score, more than 30% higher than the second-best SLVAE. Although we have a lower recall rate, a decrease of 0.09 only indicates around 1 node is not recalled from the ground truth, as only around 10 nodes are chosen as the ground truth sources in each infection. However, an increase of 0.3 in precision represents around 6 nodes correctly identified without false positives. Therefore, precision should be considered the more critical metric in source localization problems than recall when the F1 scores are similar, and SIDSL demonstrates the strongest competitiveness among the four methods.
Appendix D Dataset Description
The detailed description of the adopted datasets is presented as follows.
Table 6. Overview of networks in the datasets. ”CC” denotes ”clustering coefficient”
Dataset Nodes Edges Mean Degree CC
@@ Appendix D Dataset Description
Net 1589 2742 1.72 0.6377
Power 4941 6594 1.33 0.0801
D.1. Real-world Propagation Dataset: Digg
The datasets selected, Digg and Twitter, represent real-world social networks where information propagation can be authentically traced, and are commonly used for evaluation in previous works (Ling et al., 2022; Huang et al., 2023b). Both include propagation cascades demonstrating the time stamps and the information diffusion trace among users of each post or message. A connection network of all users is also provided in each dataset. Both datasets are pertinent to our study because they exemplify real-world dynamics of information spread.
Digg (Rossi and Ahmed, 2015) is real-world social network data showcasing voting records of stories that made it to Digg’s front page in 2009, with each story’s spread counted as one diffusion cascade. We randomly choose 100 stories to form our dataset. The nodes (voters) involved in these stories form a subgraph of the original one, where the links represent the friendship of voters. The statistics of this friendship network are shown in Table 6. Drawing an analogy to the spread of a virus during a pandemic, it is often difficult to detect the virus at the very beginning, but after some time has passed—such as when the manifestation of symptoms—we can observe the infection status of the population. As a result, for each story cascade, we choose the top 10% of nodes and 30% of nodes as diffusion sources and observed influenced nodes based on their influenced time.
D.2. Real-world Propagation Dataset: Twitter
The Twitter (Yang et al., 2021) dataset is a collection of social network and public tweets written in English that were posted on the social media platform Twitter (a.k.a X) from March 24th to April 25th, 2012. The network statistics are shown in Table 6. Each tweet can be counted as one propagation cascade. Same as Digg, for each cascade, we choose the top 10% of nodes and 30% of nodes as diffusion sources and observed influenced nodes based on their influenced time.
D.3. Community Q&A Dataset: Android and Christianity
In Stack Exchange Q&A communities, knowledge dissemination occurs through a series of social interactions. When users post questions, provide answers, leave comments, or cast votes, these activities form a complex social network structure. Each post is marked with specific tags (such as ”google-pixel-2” in the Android community), and these posts with identical tags, arranged chronologically, create what we call ”cascades.” Our study focuses on two Stack Exchange communities with distinctly different themes: the Android community, which specializes in mobile device technical support, and the Christianity community, which centers on religious and theological discussions.
Same as Digg, for each cascade, we choose the top 10% of nodes and 30% of nodes as diffusion sources and observed influenced nodes based on their influenced time.
D.4. Synthetic Dataset
We synthesize propagation data under SIR, IC, and LT models on these three real-world networks: jazz, network science and power grid. These networks differ in scale, sparsity, and clustering characteristics, which enables us to investigate the model’s performance on different types of networks. The statistic overview is presented in Table 6. For the propagation models, the propagation properties of the SIS infection model are determined by the inherent characteristics of the disease, applying homogeneity for all nodes/edges, i.e. the infection and recovery rates in SIR are constant; for the IC and the LT influence model, the heterogeneous propagation probability of each edge is considered, which is set to be inversely proportional to the degree of the target node. This aligns with real-world propagation patterns, where nodes with more connections tend to be less receptive to information from each neighbor.
  • Jazz (Rossi and Ahmed, 2015). The provided dataset is a network of collaborations among Jazz musicians. Each node in the network represents a musician, and every edge connects two musicians who have performed together in a band. Rumors or infectious diseases are applicable to be propagated on such networks. We randomly choose 5% of nodes to be the spreading sources of each propagation and use SIS, IC, or LT models to simulate 100 steps or simulate until convergence.
  • Network Science (Net) (Rossi and Ahmed, 2015). This is a coauthorship network of scientists working on network theory. Nodes represent scientists and edges represent collaborations. Influential information can be propagated on such networks. We randomly choose 0.5% of nodes to be the spreading sources of each propagation and use SIS, IC, or LT models to simulate 100 steps or simulate until convergence.
  • Power Grid (Power) (Watts and Strogatz, 1998). This is a topology network of the power grid network across the Western United States. In this network, each connection denotes a transmission line for electrical power. The nodes signify one of three components: a power generation unit, a transformer, or a distribution substation. Blackouts can be propagated on such a network. We randomly choose 0.5% of nodes to be the spreading sources of each propagation and use SIS, IC, or LT models to simulate 100 steps or simulate until convergence.
Appendix E Baselines
We compare the performance of SIDSL against three state-of-the-art baselines of source localization methods using propagation snapshot observations. To the best of our knowledge, these methods are the only ones that illustrate their superiority against other works on locating sources without knowing the underlying propagation pattern, which is the same as ours. The detailed information is presented as follows.
  • NetSleuth (Prakash et al., 2012) utilizes a minimum description length approach to filter nodes from multiple sources, yet it is exclusively designed to operate within the Susceptible-Infected (SI) model framework.
  • LPSI (Wang et al., 2017) is a novel method for detecting multiple sources of information diffusion in networks without a predefined propagation model, leveraging the concept of source prominence and label propagation to identify probable sources based on local peaks in the propagation landscape. In our experiment, the parameter $\alpha$ in LPSI is determined by testing it among the values {0.1, 0.3, 0.5, 0.7, 0.9} for each evaluation dataset and then selecting the best one.
  • GCNSI (Dong et al., 2019) introduces a deep learning approach for identifying multiple rumor sources in social networks without needing the underlying propagation model, using graph convolutional networks to enhance prediction precision through spectral domain convolution and multi-order neighbor information. The setting of this model follows the description in (Dong et al., 2019).
  • SLVAE (Ling et al., 2022) is a probabilistic framework designed to tackle the challenge of source localization in graph diffusion problems using a variational autoencoder approach to quantify uncertainty and leverage prior knowledge. We follow the original implementation in the paper, tune the learning rate from 0.001 to 0.05, and select the best one.
  • TGASI (Hou et al., 2023) is a sequence-to-sequence framework for multiple rumor source detection that considers heterogeneous user behavior in time-varying scenarios. It uses a GNN-based encoder to generate multiple features and a GRU-based decoder with temporal attention to infer sources. TGASI is designed with transferability and uses a unique loss function.
  • DDMSL (Yan et al., 2024) proposes a novel probabilistic model for source localization and diffusion path reconstruction in complex networks. By formulating information propagation as a discrete diffusion process, DDMSL employs a reversible residual network to construct a denoising-diffusion model in discrete space. This approach allows for both accurate source identification and comprehensive reconstruction of information diffusion pathways.
Appendix F Experiments and Implementation Details
For each dataset, the ratio of training, validation, and testing portion is 6:1:1. For the diffusion framework of SIDSL, we use $T=500$ maximum diffusion timestep and linear schedule for noise scheduling. In the denoising network, we leverage a 2-layer graph convolutional network (GCN) to forward the LPSI estimation $X_{est}$. The GNN encoder and decoder comprise 3-layer GCNs with a hidden dimension of 128. The residual GNN of the conditioning module is a 2-layer GCN, with a hidden dimension of 8. The learning rate is searched from 0.01, 0.005, 0.001, and the maximum number of training epochs is set to 500 for all datasets. In the few-shot learning experiments, the maximum pretrain/finetune epoch is set to 300. We train our model using Adam optimizer and a learning rate scheduler with a linear decay. Our model is trained on a single NVIDIA GeForce RTX 2080 Ti. The code implementation can be found at https://anonymous.4open.science/r/ASLDiff-4FE0.
Appendix G Limitations
Our proposed method also exhibits certain limitations. Our approach may, to some extent, depend on the accuracy of the advice provided by soft labels, despite our application of various sophisticated designs to enhance the model’s adaptability. As a result, when confronted with more complex scenarios, our method might reveal limitations. On the other hand, the sampling speed of our multi-step diffusion model may be slower compared to some deep learning methods, which could become a bottleneck for applications requiring real-time localization. While computational constraints currently limit our model’s direct application to million-node networks, the core principles we developed can be integrated into hierarchical approaches. This hierarchical strategy would effectively reduce the network scale, allowing us to leverage our method’s proven strength in accurate source localization for moderately-sized networks. We will continue to conduct in-depth research in these areas.

### v3

@@ 5.4. Analysis of Pretraining on Simulation Data
To investigate how pretraining on simulation data affects the model’s few-shot/zero-shot performance under different data volumes, we evaluate SIDSL’s F1 scores across the pretraining dataset with different propagation model combinations (IC, LT, and IC+LT with a ratio of 1:1) and compare them with non-pretrained (NP) SIDSL. We then test them under different ratios of the finetune dataset of real-world propagation data. The results of four datasets are shown in Figure 4.
Finetune data volume analysis. Across all four datasets, models with pretraining consistently outperformed non-pretrained (NP) models in low-data scenarios. Under the IC+LT combination, the pretrained models achieved optimal F1 scores on Digg and Twitter using just 5% of training data, while NP models required 50%. Similarly, on Android and Christianity datasets, pretrained models reached near-peak performance (within 0.033 and 0.018 F1) with only 20% of the data, compared to NP models, which needed 50%. These findings suggest that aligning pretraining patterns with specific propagation characteristics benefits model performance.
Pattern analysis. On the Digg dataset, SIDSL pretrained with IC+LT and IC configurations outperform those using LT, indicating that IC patterns better align with Digg’s propagation characteristics, as also shown in (Zhang et al., 2024). For Android and Christianity Q&A datasets, IC+LT and LT configurations performed better in few-shot scenarios. This is likely because user participation in these platforms’ information cascades (e.g., comments) is driven by cumulative influence from multiple sources (LT) rather than single independent triggers (IC). On Twitter, the IC+LT configuration performs the best, demonstrating that pattern diversity enhances pretraining effectiveness on this dataset.
@@ 5.5. Ablation Study
We conduct ablation studies to investigate the significance of key components in SIDSL. For the first ablation model, SIDSL w/o SBD (structure-prior biased diffusion), instead of computing estimation through Equation 1, we directly substitute it with the node having the *highest closeness centrality* in the infection subgraph. For the second ablation, SIDSL w/o GNN-LP, we replace the observation state encoding with a standalone GNN while maintaining an equivalent parameter count. Table 4 presents the ablation results on Digg and Christianity datasets under both full (100%) and limited (10%) training data conditions. Removing either component leads to decreased overall performance under both conditions. There are two more findings: First, removing SBD leads to more severe performance degradation compared to removing GNN-LP, with the impact most pronounced in recall metrics. The substantial decline in recall (TP/(TP+FN)) indicates reduced true positive predictions. The decline in recall exceeds that of precision (TP/(TP+FP)), implying fewer positive predictions overall (TP+FP). This demonstrates SBD’s role in maintaining predictive balance by ensuring adequate positive predictions. Second, the impact of removing GNN-LP shows greater impact under limited data (F1 drops: -17% to -21% at 10% vs. -4% to -7% at full data), demonstrating GNN-LP’s effectiveness in low-resource scenarios.
@@ 5.6. Time Cost Analysis
In inference time, our model is the second-highest because of the iterative nature of the DDPM-based denoising process. We opt for DDPM as our foundation due to its classical design and proven effectiveness. Note that SIDSL’s architecture is fully compatible with more computationally efficient diffusion variants, such as DDIM (Song et al., 2020), which could substantially reduce the current computational overhead. This flexibility, combined with our model’s transfer capabilities, makes SIDSL particularly resource-efficient in practical deployments.
@@
5.7. Results on Synthetic Data
@@
It is a common practice in previous works (Ling et al., 2022; Yan et al., 2024; Wang et al., 2022) to evaluate methods’ performance on synthetic data with established propagation patterns. Following (Ling et al., 2022; Wang et al., 2022), we conducted experiments on two commonly evaluated networks: Net Science (Net) and Power Grid(Power). Using both IC and LT models, we simulated 100 steps until convergence to generate two sets of synthetic data. Strong baseline models were trained and tested on these datasets, with results shown in Figure 6. Our method outperforms the baselines, achieving F1 score improvements of 7.6% and 4.7% under the IC model, and 2.7% and 1.7% under the LT model on the two datasets, respectively. These results demonstrate our method’s ability to accurately identify sources with established propagation patterns.
6. Conclusion and Future Work
We introduced SIDSL, a structure-aware diffusion model that significantly enhances source localization in sparse graphs by integrating structural priors. Our framework demonstrates a notable improvement in F1 scores by 7.5-13.3% over leading-edge methods, proving especially robust in few-shot and zero-shot scenarios where it maintains high performance with as little as 10% of training data. This highlights its strong generalization capabilities and practical value in data-scarce environments. However, the iterative nature of the diffusion process introduces computational overhead, which may limit its use in real-time applications. While the current architecture is not suited for million-node networks, its core principles can be adapted for hierarchical frameworks to analyze large-scale graphs. Future work will focus on improving computational efficiency and extending the framework to handle a wider variety of source types. These limitations also open up avenues for further research to enhance the model’s scalability and robustness.
Appendix
Table 5. Overview of networks. $C$: clustering coefficient.
Dataset Nodes Edges Mean Degree $C$
@@ 6. Conclusion and Future Work
Net Science 1589 2742 1.72 0.6377
Power Grid 4941 6594 1.33 0.0801
Appendix A Dataset Description
The study utilizes both real-world and synthetically generated propagation data. This section details the datasets employed in our study, with network statistics summarized in Table 5.
The cascades in real-world propagation datasets represent information diffusion. These include Digg(Rossi and Ahmed, 2015), with cascades formed by votes on stories; Twitter(Yang et al., 2021), where each tweet’s spread is a cascade; and the Stack Exchange communities Android and Christianity (Huang et al., 2023b), where cascades are chronologically ordered posts with identical tags. The top 10% of nodes ranked by influence time were designated as diffusion sources for each cascade, with the top 30% considered as observed influenced nodes.
Synthetic propagation datasets were generated using SIR, IC, and LT models on three real-world networks: Jazz(Rossi and Ahmed, 2015), Net Science(Rossi and Ahmed, 2015), and Power Grid (Watts and Strogatz, 1998). The SIR model used constant infection and recovery rates. For IC and LT models, edge propagation probabilities were inversely proportional to the target node’s degree. Simulations ran for 100 steps or until convergence. The Jazz network represents musician collaborations; 5% of nodes were randomly selected as sources. The Net Science network is a coauthorship graph of network scientists,and the Power Grid network maps the Western US electrical grid topology; 0.5% of nodes were randomly selected as sources in both datasets.
Table 6. Comparison of localization methods.Obs.: Observation input (single/multiple snapshots); PI: Prior-knowledge informed; Patterns: Applicable diffusion patterns.
Category Method Indeterminacy Patterns Obs. PI
Rule-based NetSleuth (Prakash et al., 2012) SI single -
OJC (Zhu et al., 2017) SI, SIR single -
LPSI (Wang et al., 2017) any single -
Data-driven GCNSI (Dong et al., 2019) any single
IVGD (Wang et al., 2022) IC single
SLVAE (Ling et al., 2022) any single
SLDiff (Huang et al., 2023b) any multiple
TGASI (Hou et al., 2023) any multiple
DDMSL (Yan et al., 2024) any multiple
PGSL (Xu et al., 2024) any single
GINSD (Cheng et al., 2024) any single
NFSL (Hou et al., 2024) any single
CRSLL (Hou et al., [n. d.]) any single
SIDSL (Ours) any single
Appendix B Comparison of Multiple Source Localization Methods
Table 6 compares mainstream source localization methods on functionality, requirements, and applications. Our method is more functional, versatile, requires less input, and addresses existing limitations, offering significant practical value. Unlike other diffusion-based methods (DDMSL, TGASI), ours doesn’t require propagation data or model parameters pre-localization. PGSL, using a flow-model, is surpassed by our diffusion model’s superior distribution modeling. GINSD addresses incomplete data (not our focus), otherwise reducing to a GAT-based baseline like GCNSI. Cross-platform works (Wang et al., 2024; Ling et al., 2024) are orthogonal and not discussed.
Appendix C Baselines
We compare SIDSL against state-of-the-art baselines that locate sources from propagation snapshots without requiring knowledge of the underlying propagation pattern.
NetSleuth (Prakash et al., 2012) employs a Minimum Description Length approach for multiple sources, specifically within the Susceptible-Infected (SI) model. LPSI (Wang et al., 2017) detects multiple sources using source prominence and label propagation; its parameter $\alpha$ was tuned from {0.1, 0.3, 0.5, 0.7, 0.9} for each dataset. GCNSI (Dong et al., 2019) utilizes Graph Convolutional Networks for multiple rumor source identification, with settings as per the original paper. SLVAE (Ling et al., 2022) is a probabilistic Variational Autoencoder for source localization that quantifies uncertainty; we used the original implementation and tuned the learning rate (0.001-0.05). TGASI (Hou et al., 2023) offers a sequence-to-sequence framework (GNN encoder, GRU decoder with temporal attention) for multiple rumor source detection, considering heterogeneous user behavior and time-variance, designed for transferability with a unique loss function. DDMSL (Yan et al., 2024) proposes a probabilistic model for source localization and diffusion path reconstruction via a discrete denoising-diffusion model with a reversible residual network.
Appendix D Experiments and Implementation Details
We partitioned datasets using a 6:1:1 train/validation/test split. SIDSL employed T=500 diffusion timesteps with a linear noise schedule, without classifier-free sampling. Grid search was conducted over encoder/decoder layers {2, 4, 6} and hidden dimensions {32, 64, 128, 256}. After comparing GNN variants (GCN, GIN, GAT, MLP), we selected a 2-layer GCN for LPSI estimation ($X_{est}$) as the denoising network. The encoder/decoder architecture consisted of 3-layer GCNs with 128-dimensional hidden layers, while the conditioning module employed a 2-layer GCN with 8-dimensional hidden layers. Models were trained using Adam optimizer with learning rates selected from {0.01, 0.005, 0.001} and linear decay, running for 500 epochs (300 for few-shot scenarios) on an RTX 2080 Ti.
Appendix E Analyzing Source Centrality in Empirical Data
Figure 7. Normalized closeness centrality (CC) on Digg. (a)(b) $C$ distributions for infected (blue) vs. source (red) nodes: (a) Density plot, with mean $C$ (dashed lines). (b) Frequency plot, orange box highlights $C$ ¿ 0.8. (c) CC probability density function for ground truth (blue) vs. predicted (red) sources.
The assumption of source centrality is validated by our empirical analysis of the Digg dataset. Figure 7(a) shows the mean normalized closeness centrality (CC) of source nodes is markedly higher than other infected nodes. Furthermore, Figure 7(b) reveals that sources constitute over 63% of nodes with CC scores above 0.8, highlighting their key structural role.
Our proposed SIDSL model is informed by this centrality principle but uses a Graph Neural Network (GNN) for adaptive learning. As shown in Figure 7(c), the CC probability density of SIDSL’s predicted sources on Digg (mean 0.7020, std 0.1444) closely matches that of the ground truth sources (mean 0.7044, std 0.1567). This statistical alignment shows our model captures empirical source distributions without introducing bias, demonstrating a robust synthesis of knowledge-guided inference and data-driven learning.
Acknowledgements.
This work was supported by the Shenzhen Ubiquitous Data Enabling Key Lab under grant ZDSYS20220527171406015, the Tsinghua Shenzhen International Graduate School-Shenzhen Pengrui Endowed Professorship Scheme of the Shenzhen Pengrui Foundation and the National Natural Science Foundation of China under U23B2030.
GenAI Usage Disclosure
The authors used Claude (Anthropic) to assist with specific formatting tasks and language editing in this work. Specifically, Claude was used to:
  • Convert pre-existing tabular data into LaTeX table format for presentation purposes
  • Provide grammar correction and language polishing of author-written text
No content generation, analysis, interpretation, or creation of new ideas was performed using generative AI tools. All original research, methodology, results, and conclusions presented in this work are entirely the authors’ own. The use of Claude was limited to technical formatting assistance and language refinement, similar to using grammar checking tools or word processing software.